OVHCloud Bare Metal Cloud Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#9926 - HG - 178.33.62.0/24
Incident Report for Bare Metal Cloud
Resolved
There is some instability on the HG network switch.

We are investigating

Update(s):

Date: 2013-12-23 16:37:43 UTC
Done

We have corrected the problem

If it becomes unstable again we will change the swithch

Date: 2013-12-23 13:24:54 UTC
We suspect there's a problem with Asic on the N5.

We're switching the traffic and will then reboot the switch with the problem.
Posted Dec 23, 2013 - 10:58 UTC